# Dynamical cluster-based optimization of tensor network algorithms for quantum circuit simulations

### v1

@@ IV.2.1 Clifford random-structure quantum circuits
Fig. 8 summarizes the results of our benchmarks for Clifford random-structure quantum circuits: the 3D plot in Fig. 8a shows, in logarithmic scale, the runtime of cluster-TEBD and standard TEBD for bond dimension $4096$, which assures $\mathcal{F}\simeq 1$ for all circuits. The plot clearly shows that the runtime of cluster-TEBD is constantly lower than that of standard TEBD, meaning that, for the considered circuit types and sizes, cluster-TEBD offers a considerable speedup over standard TEBD. If we quantify this speedup by computing the ratio between standard and cluster-TEBD runtimes, we find out that cluster-TEBD is at least $2.276$ times faster than standard TEBD, with a peak speedup of more than $6$ times in the case of $40$ qubit, $40$ layer Clifford random-structure quantum circuits. Figs. 8b-8c are insets of the 3D plot, with the number of qubits fixed at 1000 and the number of layers fixed at 100, respectively.
@@
Fig. 8d shows the fidelity of the final states obtained from the standard TEBD and cluster-TEBD algorithms when simulating Clifford random-structure quantum circuits with $100$ layers, plotted as a function of the number of qubits in each circuit, for different values of $\chi_{\mathrm{max}}$. It is possible to notice that using $\chi_{\mathrm{max}}=4096$ as maximum bond dimension guarantees fidelity $\mathcal{F}\simeq 1$ even for the largest circuit analyzed in this work. Notably, using $\chi_{\mathrm{max}}=128$ produces a fidelity close to 00 for all Clifford random-structure quantum circuits with $100$ gate layers, while $\chi_{\mathrm{max}}=512$ and $\chi_{\mathrm{max}}=1024$ only get slightly close to fidelity $1$ for circuits with tens of qubits. Additionally, while the fidelities for standard TEBD and cluster-TEBD are around the same order of magnitude in every point, we observe an improvement in the fidelity for cluster-TEBD simulations, which may be attributed to the larger number of layers simulated exactly, without any SVD or truncation step.
IV.2.2 Non-Clifford random-structure quantum circuits
@@
Figure 9: Comparison of cluster-TEBD and TEBD in simulations of non-Clifford random-structure quantum circuits. (a) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of non-Clifford random-structure quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$ (b) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of $1000$-qubit non-Clifford random-structure quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$ and varying number of layers. (c) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD exact simulations of $80$-layer non-Clifford random-structure quantum circuits with bond dimension $4096$ and varying number of qubits. (d) Fidelity of standard TEBD and cluster-TEBD simulations of $80$-layer non-Clifford random-structure quantum circuits with varying number of qubits.
@@
We now benchmark random-structure quantum circuits again, this time with a $50\%$ probability of introducing non-Clifford gates during the evolution, to test the performance of cluster-TEBD on non-stabilizer circuits. We use the same number of qubits and bond dimensions as in the Clifford case, stopping at a circuit depth of $80$ layers. This is because, even with the largest bond dimension used in our simulations, the final fidelity for non-Clifford random-structure quantum circuits with more than $80$ gate layers turns out to be relatively low for both standard and cluster-TEBD. However, a maximum bond dimension $\chi_{\mathrm{max}}=4096$ can still accurately represent the final state of all analyzed circuits of this kind with up to $80$ layers, achieving a fidelity of $\mathcal{F}\simeq 1$.
@@
Fig. 9 summarizes the results of our benchmarks, where we average out the runtime and fidelity results of $20$ different samples. Looking at the 3D plot in Fig. 9a, showing runtimes, in logarithmic scale, for simulations with $\chi_{\mathrm{max}}=4096$, it is evident that the cluster-TEBD runtime is again constantly lower than that of standard TEBD for the considered circuit sizes. We observe that non-Clifford gates increase entanglement, and consequently the bond dimension, more rapidly than Clifford gates. As a result, the number of clustered qubits decreases more quickly compared to stabilizer circuits, reducing the effectiveness of cluster-TEBD slightly. Nevertheless, cluster-TEBD still presents a speedup compared to standard TEBD, going, on average, from $1.5$, for circuits with $1000$ qubits and $80$ layers, up to $5.5$, for circuits with $40$ qubits and $40$ layers.
@@
The increased simulation difficulty of the non-stabilizer version of random-structure quantum circuits is confirmed by the fidelity plot in Fig. 9d. While representing the final state of these non-stabilizer circuits with intermediate bond dimensions is evidently much harder than in the stabilizer case, $\chi_{\mathrm{max}}=4096$ again allows us to simulate them exactly. Another interesting feature to note is that the cluster-TEBD advantage in improving final fidelity for low bond dimensions is even more pronounced with non-Clifford gates compared to the Clifford case. This may be due to the larger number of non-Clifford gates being contracted exactly in the cluster-TEBD procedure.
@@ IV.2.3 Shor’s algorithm
Figure 10: Runtime of standard TEBD and cluster-TEBD exact simulations of Shor’s algorithm for factoring $15$, $21$ and $35$.
The cluster-TEBD algorithm demonstrates significant advantages when applied to the simulation of random-structure quantum circuits, even as the bond dimension quickly becomes very large. In the following, we assess its performance in simulating a representative example of quantum algorithms, where we anticipate that the bond dimension will remain more constrained throughout the evolution [58, 31]. In this scenario, we expect to group large clusters of entanglement throughout the whole simulation process, leading to improved overall performance.
Shor’s algorithm for factoring constitutes a great example of such a task, due to the sparse presence of gates in each layer. Even though the depth of these circuits is of the order of tens, or even hundreds of thousands gate layers, each of these layers contains a limited amount of gates. This is because of the structure of the algorithm itself, which is mainly composed of rotations and $CNOT$ gates on consecutive qubit pairs. Additionally, the presence of long range interactions forces the application of consecutive $SWAP$ gates, each belonging to a single layer, to operate on distant qubits.
We simulate Shor’s algorithm for factoring the numbers $15,21,35$, using precision values of $\epsilon~{}=~{}10^{-3}$, $\epsilon~{}=~{}10^{-4}$, and $\epsilon~{}=~{}10^{-5}$, respectively. Following Beauregard’s implementation of Shor’s algorithm [56], the quantum circuits for factoring these numbers display the following characteristics:
  • For $N=15$ with $\epsilon=10^{-3}$: The circuit requires $29$ qubits and $70122$ gates, arranged in $49912$ layers.
  • For $N=21$ with $\epsilon=10^{-4}$: The circuit requires $37$ qubits and $158128$ gates, arranged in $116833$ layers.
  • For $N=35$ with $\epsilon=10^{-5}$: The circuit requires $44$ qubits and $299193$ gates, arranged in $229594$ layers.
Fig. 10 shows the runtime of cluster-TEBD and standard TEBD for simulating this quantum circuit implementation of Shor’s algorithm for the specified numbers. The cluster-TEBD runtime is significantly lower than standard TEBD’s, as in the case of random-structure quantum circuits, but the speedup in this case is much higher and more consistent across the analyzed factorization problems. In numbers, cluster-TEBD is more than $8$ times faster than standard TEBD at factoring $15$ with Shor, and around $5.5$ times faster at factoring $21$ and $35$.
@@ IV.3 DMRG with higher number of layers
Thanks to the adaptive clustering algorithm introduced in Sec. III.2, which generates the grouped MPS according to the circuit structure, it becomes possible to employ the DMRG algorithm to simulate quantum circuits with irregular structures. This allows us to run benchmarks for DMRG on random-structure quantum circuits. In particular, we test the non-stabilizer variation, choosing sizes of $60$, $80$, $100$ qubits, and $60$, $80$ layers, with the goal of understanding how many layers can be exactly contracted in a single step. While in the original work by Ayral et al. [30] the largest number of layers $L_{\mathrm{max}}$ in each iteration employed was $4$, we vary $L_{\mathrm{max}}$ from $2$ to $20$.
@@
Figure 11: Results of DMRG benchmarks on 20 realizations of non-Clifford random-structure quantum circuits of various sizes with maximum bond dimensions $\chi_{\mathrm{max}}^{\mathrm{DMRG}}=256$ and $\chi_{\mathrm{max}}^{\mathrm{SVD}}=4096$, and $n_{s}=3$ maximum sweeps. Top: Average runtime of DMRG simulations as a function of the maximum number of layers per step $L_{\mathrm{max}}$. Bottom: Average fidelity of DMRG simulations as a function of the maximum number of layers per step $L_{\mathrm{max}}$.
In the bottom plot from Fig. 11, we show the evolution of the average fidelity across the simulated circuit sizes as a function of $L_{\mathrm{max}}$, demonstrating an advantage for larger $L_{\mathrm{max}}$ values. This improvement in fidelity is due to several reasons. First, the adaptive grouping procedure requires the grouped MPS to be converted back into a standard MPS with successive SVDs. Setting a lower $L_{\mathrm{max}}$ increases the frequency of these decompositions throughout the algorithm. Since each decomposition step can introduce fidelity loss through truncation, a higher $L_{\mathrm{max}}$ helps to reduce these potential reductions. Additionally, for growing $L_{\mathrm{max}}$, the number of gate layers contracted exactly within a single iteration rises accordingly.
Appendix B provides additional results for the DMRG algorithm, where we compare the runtime performance of using the adaptive grouping protocol introduced in this work against a fixed grouping.
@@ V Discussion
In this work, we have optimized tensor network algorithms for simulating quantum circuits with finite fidelity by implementing entanglement clustering protocols. We designed a variation of the TEBD algorithm, called the “cluster-TEBD” algorithm, which allows the exact contraction of multiple layers of quantum gates in a single step. This algorithm employs an adaptive clustering technique that groups gates in different layers based on the arrangement of entangling operations, increasing each cluster’s size until the selected layers can be contracted exactly. Moreover, we implemented an automatic clustering scheme for the DMRG algorithm, using hypergraph partitioning [35], where the qubit grouping adaptively changes in each iteration according to the arrangement of entangling gates in the simulated circuit portion. These enhancements allowed efficient simulations of quantum circuits where entangling operations are scattered across a large chunk of gate layers, rather than condensed into few layers.
Our results demonstrate that cluster-TEBD significantly outperforms standard TEBD in both runtime and fidelity when simulating quantum circuits with random gates and structure. Furthermore, we have shown that this advantage is not only maintained but also considerably enhanced when applied to the simulation of a more practical task, specifically Shor’s quantum algorithm for factoring. Additionally, using the adaptive clustering algorithm in DMRG, we have focused on studying the effect of the choice of the maximum number of contracted layers per iteration $L_{\mathrm{max}}$. We have demonstrated that, for a given quantum circuit to simulate, there exists a range of valid values for $L_{\mathrm{max}}$, even larger than what was considered in previous works [30]. Moreover, we have verified that, using the adaptive grouping technique, the final fidelity of the simulation increases with $L_{\mathrm{max}}$.
The framework introduced in this work could be adapted to other tensor network ansätze beyond MPS. For instance, it could be extended to tree tensor networks [61, 62] and hybrid quantum-classical structures [63, 64], combining them with adaptive techniques to build the network according to the structure of the evolution [65, 66]. With this modification, it would be possible to extend our numerical methods to quantum circuits of 2D qubit topologies.
Another promising aspect is the potential for parallelization in cluster-TEBD, as the contraction of each cluster can be performed independently before orthogonalizing the state to measure observables. Introducing large-scale parallelism, such as MPI-based methods, could further enhance the performance [67].
@@ Acknowledgments
The authors acknowledge financial support from the following institutions: the European Union (EU), via the NextGenerationEU projects “CN00000013 - Italian Research Center on HPC, Big Data and Quantum Computing (ICSC)” and “PE0000023 - National Quantum Science and Technology Institute (NQSTI)”; the Italian Ministry of University and Research (MUR), via PRIN2022 project “2022NZP4T3 - QUEXO” (CUP D53D23002850006), and via the Department of Excellence grants 2023-2027 project “Quantum Sensing and Modelling for One-Health (QuaSiModO)”; the Italian Istituto Nazionale di Fisica Nucleare (INFN), via the projects QUANTUM and NPQCD; the University of Bari, via the 2023-UNBACLE-0244025 grant. This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. We acknowledge computational resources by CINECA and the University of Bari and INFN cluster RECAS [68]. The numerical simulations were performed using the following libraries: ITensors.jl [69], for management of tensors and MPS, opt_einsum [36] and cotengra [21], for heuristics and optimized exact contractions.
Appendix A Gates used in random-structure quantum circuits
In this Appendix, we define the gates used in random-structure quantum circuits. Let us start with the Clifford single-qubit gates, which are simply

### v2

@@ IV.2.1 Clifford random-structured quantum circuits
Fig. 8 summarizes the results of our benchmarks for Clifford random-structured quantum circuits. The three-dimensional (3D) plot in Fig. 8(a) shows, in logarithmic scale, the runtime of cluster-TEBD and standard TEBD for bond dimension $4096$, which assures $\mathcal{F}\simeq 1$ for all circuits. The plot clearly shows that the runtime of cluster-TEBD is constantly lower than that of standard TEBD, meaning that, for the considered circuit types and sizes, cluster-TEBD offers a considerable speedup over standard TEBD. If we quantify this speedup by computing the ratio between standard and cluster-TEBD runtimes, we find out that cluster-TEBD is at least $2.276$ times faster than standard TEBD, with a peak speedup of more than $6$ times in the case of $40$ qubit, $40$ layer Clifford random-structured quantum circuits. Figs. 8(b)-8(c) are insets of the 3D plot, with the number of qubits fixed at 1000 and the number of layers fixed at 100, respectively.
@@
Fig. 8(d) shows the fidelity of the final states obtained from the standard TEBD and cluster-TEBD algorithms when simulating Clifford random-structured quantum circuits with $100$ layers, plotted as a function of the number of qubits in each circuit, for different values of $\chi_{\mathrm{max}}$. It is possible to notice that using $\chi_{\mathrm{max}}=4096$ as maximum bond dimension guarantees fidelity $\mathcal{F}\simeq 1$ even for the largest circuit analyzed in this work. Notably, using $\chi_{\mathrm{max}}=128$ produces a fidelity close to 0 for all Clifford random-structured quantum circuits with $100$ gate layers, while $\chi_{\mathrm{max}}=512$ and $\chi_{\mathrm{max}}=1024$ only get slightly close to fidelity $1$ for circuits with tens of qubits. Additionally, while the fidelities for standard TEBD and cluster-TEBD are around the same order of magnitude at every point, we observe an improvement in the fidelity for cluster-TEBD simulations, which may be attributed to the larger number of layers simulated exactly, without any SVD or truncation step.
IV.2.2 Non-Clifford random-structured quantum circuits
@@
Figure 9: Comparison of cluster-TEBD and TEBD in simulations of non-Clifford random-structured quantum circuits. (a) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of non-Clifford random-structured quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$ (b) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD simulations of $1000$-qubit non-Clifford random-structured quantum circuits with maximum bond dimension $\chi_{\mathrm{max}}=4096$ and varying number of layers. (c) Runtime, in logarithmic scale, of standard TEBD and cluster-TEBD exact simulations of $80$-layer non-Clifford random-structured quantum circuits with bond dimension $4096$ and varying number of qubits. (d) Fidelity of standard TEBD and cluster-TEBD simulations of $80$-layer non-Clifford random-structured quantum circuits with varying number of qubits.
@@
We now benchmark random-structured quantum circuits again, this time with a $50\%$ probability of introducing non-Clifford gates during the evolution, to test the performance of cluster-TEBD on nonstabilizer circuits. We use the same number of qubits and bond dimensions as in the Clifford case, stopping at a circuit depth of $80$ layers. This is because, even with the largest bond dimension used in our simulations, the final fidelity for non-Clifford random-structured quantum circuits with more than $80$ gate layers turns out to be relatively low for both standard and cluster-TEBD. However, a maximum bond dimension $\chi_{\mathrm{max}}=4096$ can still accurately represent the final state of all analyzed circuits of this kind with up to $80$ layers, achieving a fidelity of $\mathcal{F}\simeq 1$.
@@
Fig. 9 summarizes the results of our benchmarks, where we average out the runtime and fidelity results of $20$ different samples. Looking at the 3D plot in Fig. 9(a), showing runtimes, in logarithmic scale, for simulations with $\chi_{\mathrm{max}}=4096$, it is evident that the cluster-TEBD runtime is again constantly lower than that of standard TEBD for the considered circuit sizes. We observe that non-Clifford gates increase entanglement, and consequently the bond dimension, more rapidly than Clifford gates. As a result, the number of clustered qubits decreases more quickly compared to stabilizer circuits, reducing the effectiveness of cluster-TEBD slightly. Nevertheless, cluster-TEBD still presents a speedup compared to standard TEBD, going, on average, from $1.5$, for circuits with $1000$ qubits and $80$ layers, up to $5.5$, for circuits with $40$ qubits and $40$ layers.
@@
The increased simulation difficulty of the nonstabilizer version of random-structured quantum circuits is confirmed by the fidelity plot in Fig. 9(d). While representing the final state of these nonstabilizer circuits with intermediate bond dimensions is evidently much harder than in the stabilizer case, $\chi_{\mathrm{max}}=4096$ again allows us to simulate them exactly. Another interesting feature to note is that the cluster-TEBD advantage in improving final fidelity for low bond dimensions is even more pronounced with non-Clifford gates compared to the Clifford case. This may be due to the larger number of non-Clifford gates being contracted exactly in the cluster-TEBD procedure.
@@ IV.2.3 Shor’s algorithm
Figure 10: Runtime of standard TEBD and cluster-TEBD simulations of Shor’s algorithm for factoring numbers of increasing bit length, from $15$ (four-bit number) to $221$ (eight-bit number). The standard TEBD runtime provided for $N=221$ is an extrapolated estimate, as the simulation reached the maximum wall time of the supercomputer; the estimate is based on the observed progress prior to termination.
The cluster-TEBD algorithm demonstrates significant advantages when applied to the simulation of random-structured quantum circuits, even as the bond dimension quickly becomes very large. In the following, we assess its performance in simulating a representative example of quantum algorithms, where we anticipate that the bond dimension will remain more constrained throughout the evolution [Chen2023QFTSmallEntanglement, Niedermeier2024QalgsTN]. In this scenario, we expect to group large clusters of entanglement throughout the whole simulation process, leading to improved overall performance.
Shor’s algorithm for factoring constitutes a great example of such a task, due to the sparse presence of gates in each layer. Even though the depth of these circuits is of the order of tens, or even hundreds of thousands gate layers, each of these layers contains a limited amount of gates. This is because of the structure of the algorithm itself, which is mainly composed of rotations and $CNOT$ gates on consecutive qubit pairs. Additionally, the presence of long range interactions poses a challenge that can be solved in two possible ways. The first approach is to replace the gate with a multisite MPO, assuming an efficient MPO representation of the gate exists. The second approach – which we adopt in this work – involves the application of consecutive $SWAP$ gates, each belonging to a single layer, to operate on distant qubits.
We simulate Shor’s algorithm for factoring the numbers $15,21,35,91,221$, with adequate precision values. Following Beauregard’s implementation of Shor’s algorithm [Beauregard2003ShorCircuit], the quantum circuits for factoring these numbers display the following characteristics:
  • For $N=15$ with $\epsilon=10^{-3}$: The circuit requires $29$ qubits and 65 59465\,594 gates, arranged in 45 42045\,420 layers.
  • For $N=21$ with $\epsilon=10^{-4}$: The circuit requires $37$ qubits and 142 106142\,106 gates, arranged in 100 847100\,847 layers.
  • For $N=35$ with $\epsilon=10^{-5}$: The circuit requires $44$ qubits and 257 841257\,841 gates, arranged in 188 282188\,282 layers.
  • For $N=91$ with $\epsilon=10^{-5}$: The circuit requires $48$ qubits and 388 432388\,432 gates, arranged in 292 635292\,635 layers.
  • For $N=221$ with $\epsilon=10^{-9}$: The circuit requires $65$ qubits and 818 434818\,434 gates, arranged in 617 553617\,553 layers.
Fig. 10 shows the runtime of cluster-TEBD and standard TEBD for simulating this quantum circuit implementation of Shor’s algorithm for the specified numbers. The cluster-TEBD runtime is significantly lower than standard TEBD’s, as in the case of random-structured quantum circuits, but the speedup in this case is much higher – cluster-TEBD is between $10$ and $20$ times faster – and more consistent across the analyzed factorization problems.
@@ IV.3 DMRG with higher number of layers
Thanks to the adaptive clustering algorithm introduced in Sec. III.2, which generates the grouped MPS according to the circuit structure, it becomes possible to employ the DMRG algorithm to simulate quantum circuits with irregular structures. This allows us to run benchmarks for DMRG on random-structured quantum circuits. In particular, we test the nonstabilizer variation, choosing sizes of $60$, $80$, $100$ qubits, and $60$, $80$ layers, with the goal of understanding how many layers can be exactly contracted in a single step. While in the original work by Ayral et al. [Ayral2023DMRG_QC] the largest number of layers $L_{\mathrm{max}}$ in each iteration employed was $4$, we vary $L_{\mathrm{max}}$ from $2$ to $20$.
@@
Figure 11: Results of DMRG benchmarks on 20 realizations of non-Clifford random-structured quantum circuits of various sizes with maximum bond dimensions $\chi_{\mathrm{max}}^{\mathrm{DMRG}}=256$ and $\chi_{\mathrm{max}}^{\mathrm{SVD}}=4096$, and $n_{s}=3$ maximum sweeps. Top: Average runtime of DMRG simulations as a function of the maximum number of layers per step $L_{\mathrm{max}}$. Bottom: Average fidelity of DMRG simulations as a function of the maximum number of layers per step $L_{\mathrm{max}}$.
In the bottom plot from Fig. 11, we show the evolution of the average fidelity across the simulated circuit sizes as a function of $L_{\mathrm{max}}$, demonstrating an advantage for larger $L_{\mathrm{max}}$ values. This improvement in fidelity is due to several reasons. First, the dynamical adaptive grouping routine requires the grouped MPS to be converted back into a standard MPS with successive SVDs. Setting a lower $L_{\mathrm{max}}$ value increases the frequency of these decompositions throughout the algorithm. Since each decomposition step can introduce fidelity loss through truncation, a higher $L_{\mathrm{max}}$ value helps to reduce these potential reductions. Additionally, for growing $L_{\mathrm{max}}$, the number of gate layers contracted exactly within a single iteration rises accordingly.
Appendix B provides additional results for the DMRG algorithm, where we compare the runtime performance of using the dynamical adaptive grouping routine introduced in this work against a fixed grouping.
@@ V Discussion
In this work, we have optimized tensor network algorithms for simulating quantum circuits with finite fidelity by implementing entanglement clustering protocols. We designed a variation of the TEBD algorithm, called the “cluster-TEBD” algorithm, which allows the exact contraction of multiple layers of quantum gates in a single step. This algorithm employs an adaptive clustering technique that groups gates in different layers based on the arrangement of entangling operations, increasing each cluster’s size until the selected layers can be contracted exactly. Moreover, we implemented an automatic clustering scheme for the DMRG algorithm, using hypergraph partitioning [Schlag2022KaHyPar], where the qubit grouping adaptively changes in each iteration according to the arrangement of entangling gates in the simulated circuit portion. These enhancements allowed efficient simulations of quantum circuits where entangling operations are scattered across a large chunk of gate layers, rather than condensed into few layers.
Our results demonstrate that cluster-TEBD significantly outperforms standard TEBD in both runtime and fidelity when simulating quantum circuits with random gates and structure. Furthermore, we have shown that this advantage is not only maintained but also considerably enhanced when applied to the simulation of a more practical task, specifically Shor’s quantum algorithm for factoring. Additionally, using the adaptive clustering algorithm in DMRG, we have focused on studying the effect of the choice of the maximum number of contracted layers per iteration $L_{\mathrm{max}}$. We have demonstrated that, for a given quantum circuit to simulate, there exists a range of valid values for $L_{\mathrm{max}}$, even larger than what was considered in previous works [Ayral2023DMRG_QC]. Moreover, we have verified that, using the dynamical adaptive grouping routine, the final fidelity of the simulation increases with $L_{\mathrm{max}}$.
The framework introduced in this work could be adapted to other tensor network ansätze beyond MPS. For instance, it could be extended to tree tensor networks [Shi2006TTN, 10.21468/SciPostPhysLectNotes.8] and hybrid quantum-classical structures [PhysRevLett.127.040501, schuhmacher2024hybridtreetensornetworks], combining them with adaptive techniques to build the network according to the structure of the evolution [Ferrari2022AdaptiveTTN, Seitz2023QC_TTN]. Additionally, our dynamical clustering methods can be used in combination with belief propagation gauging, which has recently proven extremely effective in tensor network simulations with Projected Entangled Pair States [Tindall2023BPGauging, Tindall2025_2DTNs, Rudolph2025_2DQC]. These modifications would allow to extend our numerical methods to simulate quantum circuits in 2D qubit geometries.
Another promising aspect is the potential for parallelization in cluster-TEBD, as the contraction of each cluster can be performed independently before orthogonalizing the state to measure observables. Introducing large-scale parallelism, such as MPI-based methods, could further enhance the performance [gabriel04:_open_mpi].
@@ Acknowledgments
The authors acknowledge financial support from the following institutions: the European Union (EU), via the NextGenerationEU projects “CN00000013 - Italian Research Center on HPC, Big Data and Quantum Computing (ICSC)” and “PE0000023 - National Quantum Science and Technology Institute (NQSTI)”; the European Union’s Horizon Europe research and innovation programme (Quantum Flagship) under the projects PASQuanS2 (Grant Agreement No. 101113690) and EuRyQa (Grant Agreement No. 101070144); the Italian Ministry of University and Research (MUR), via PRIN2022 project “2022NZP4T3 - QUEXO” (CUP D53D23002850006), and via the Department of Excellence grants 2023-2027 project “Quantum Sensing and Modelling for One-Health (QuaSiModO)”; the Italian Istituto Nazionale di Fisica Nucleare (INFN), via the projects QUANTUM and NPQCD; the University of Bari, via the 2023-UNBACLE-0244025 grant. This research is part of the Munich Quantum Valley, which was supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. We acknowledge computational resources by CINECA and the University of Bari and INFN cluster RECAS [ReCaS]. The numerical simulations were performed using the following libraries: ITensors.jl [Fishman2022ITensors], for management of tensors and MPS, opt_einsum [Smith2018opt_einsum], and cotengra [Gray2021HyperExact], for heuristics and optimized exact contractions.
Data availability
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A Gates used in random-structured quantum circuits
In this Appendix, we define the gates used in random-structured quantum circuits. Let us start with the Clifford single-qubit gates, which are simply